# Linear time-and-space-invariant relaxation systems

### v1

@@ I Introduction
The remainder of this paper is organized as follows. In Section II, we discuss how diffusion is a natural example of spatio-temporal relaxation. Section III contains notation and preliminaries. In Section IV, we define and characterize (internal) relaxation for LTSI systems. In Section V, we show that internal relaxation implies impedance passivity with storage that is determined from the external behaviour. In Section VI, we formalize the connection between internal and external by establishing properties of the Hankel operator of an LTSI relaxation system. Finally, Section VII summarizes our findings and discusses directions for further work.
@@ II Diffusion and relaxation
In this section, we will demonstrate that the diffusion process is a natural example of spatio-temporal relaxation. Historically, the relaxation property was first studied systematically in electrical circuits composed exclusively of one type of energy storage element—RC or RL circuits—which represent canonical examples of LTI relaxation systems [8]. These classical results naturally extend to spatially distributed systems by considering an infinite transmission line composed of resistors and capacitors periodically arranged along a spatial axis. Figure 1 illustrates the schematic of such an infinite distributed RC transmission line. Taking the spatial discretization interval $\Delta x\rightarrow 0$, this discrete electrical network is precisely governed by the one-dimensional diffusion equation [9], i.e.
@@
where $V(t,x)$ is the voltage across the capacitor and $I(t,x)$ is the applied current at time $t$ and position $x$. Introducing the variables $y(t,x)=V(t,x)$, $u(t,x)=I(t,x)/C$, and the diffussion coefficient $\alpha=1/RC$, the latter simplifies to the standard (heat) diffusion equation
@@
$y(t,x)=\int_{-\infty}^{t}\int_{-\infty}^{\infty}g(t-\tau,x-\xi)u(\tau,\xi)\,\textrm{d}\xi\textrm{d}\tau$ (3)
@@
The heat kernel clearly exhibits the relaxation property spatially. In particular, Figure 2(a) illustrates the spatial profile of (4) at various times $t$. It shows the relaxation property, as the spatial profile flattens monotonically towards equilibrium as $t$ increases. However, Figure 2(b) shows the kernel’s temporal behavior at a fixed location, highlighting that the kernel is not a completely monotone function in time, as one might expect from a classical LTI relaxation system.
@@
The diffusion process provides a clear physical connection from classical relaxation theory to its infinite-dimensional, spatially extended counterpart. This is further reinforced by the recent work [11], where it is explicitly demonstrated that the spatial discretization of the diffusion equation leads to a finite-dimensional LTI relaxation system. In this paper, we will extend the theory of LTI relaxation systems to a class of LTSI system, with the diffusion process as a canonical example of an LTSI relaxation system.
$x$$t=1$$t=2$$t=5$$t=20$
$t$$x=0$$x=0.5$$x=1$$x=2$
@@ III Notation and preliminaries
[…]
@@ III-A Square integrable functions and the Fourier transform
$\langle f,g\rangle_{L_{2}(X,\mathcal{X})}=\int_{X}\langle f(x),g(x)\rangle_{\mathcal{X}}\,\textrm{d}x$ (7)
@@
$\hat{f}(\omega)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}^{s}}e^{-j\langle\omega,x\rangle}f(x)\,\textrm{d}x,$ (8)
@@
It is well-known [12] that the Fourier transform converges for absolutely integrable functions. Due to Plancherel’s theorem, the Fourier transform can be extended to a unitary operator $\mathcal{F}\in\mathscr{B}(L_{2}(\mathbb{R}^{s},\mathbb{C}^{n}))$, i.e., $\mathcal{F}^{-1}=\mathcal{F}^{\ast}$. Since $\mathcal{F}$ agrees with the Fourier transform on the dense subspace of absolutely integrable functions, we will refer to $\mathcal{F}f$ as the Fourier transform of $f$, and we will write $\hat{f}=\mathcal{F}f$.
@@
A key property of the Fourier transform that we will exploit in this paper is that translation-invariant operators are transformed into multiplication operators. The Fourier transform of a densely defined operator $\mathcal{A}$ on $L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$ is the densely defined operator $\hat{\mathcal{A}}=\mathcal{F}\mathcal{A}\mathcal{F}^{\ast}$ on $L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$, such that $\hat{\mathcal{A}}\hat{f}\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$ is the Fourier transform of $\mathcal{A}f\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$. For each $\bar{x}\in\mathbb{R}^{s}$, the translation operator $\mathcal{T}_{\bar{x}}\in\mathscr{B}(L_{2}(\mathbb{R}^{s},\mathbb{C}^{n}))$ is given by $(\mathcal{T}_{\bar{x}}f)(x)=f(x+\bar{x})$. A densely defined operator $\mathcal{A}$ is translation-invariant if $\mathcal{T}_{\bar{x}}:\mathscr{D}(\mathcal{A})\to\mathscr{D}(\mathcal{A})$ and $\mathcal{T}_{\bar{x}}\mathcal{A}=\mathcal{A}\mathcal{T}_{\bar{x}}$ for all $\bar{x}\in\mathbb{R}^{s}$. The Fourier transform of a translation-invariant operator is a multiplication operator [12, Section 3.6], i.e., there exists a function $\hat{A}:\mathbb{R}^{s}\to\mathbb{C}^{n\times n}$ such that $(\hat{\mathcal{A}}\hat{f})(\omega)=\hat{A}(\omega)\hat{f}(\omega)$. The function $\hat{A}$ is called the symbol of the operator $\hat{\mathcal{A}}$. It is well-known that the operators $\hat{\mathcal{A}}$ and, thus, $\mathcal{A}$ are bounded if and only if the symbol $\hat{A}$ is bounded.
@@
Throughout this paper, we will consider functions that depend on a temporal variable $t$ and a spatial (frequency) variable $x$ ($\omega$). We will denote the dependence on the spatial (frequency) variable with a subscript, e.g., we will write $f_{x}$ instead of $f(x)$. In particular, we will often consider functions of the form $f:\mathbb{T}\to L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$, where $\mathbb{T}$ is a temporal domain and $\mathbb{R}^{s}$ is a spatial (frequency) domain. We interpret such functions as mapping the temporal variable $t\in\mathbb{T}$ to a spatial (frequency) profile $f(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$. We will denote the dependence on the spatial (frequency) variable with a subscript, i.e., we will write $f_{x}(t)=f(t)(x)$, so that the function $f_{x}:\mathbb{T}\to\mathbb{C}^{n}$ is the temporal trajectory at a given point in space $x\in\mathbb{R}^{s}$. Finally, we will write $\hat{f}$ to denote the spatial Fourier transform of $f$, i.e., the function $\hat{f}:\mathbb{T}\to L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ such that $\hat{f}(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ is the Fourier transform of $f(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ for all $t\in\mathbb{T}$.
@@ III-B LTI relaxation systems
$\Sigma:\left\{\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}z(t)&=Az(t)+Bu(t),\\ y(t)&=Cx(t),\end{aligned}\right.$ (9)
@@
$y(t)=\int_{-\infty}^{t}g(t-\tau)u(\tau)\,\textrm{d}\tau$ (10)
The system $\Sigma$ is of relaxation type [6] if $m=p$ and its impulse response $g$ is devoid of any oscillatory behaviour, expressed as $g$ being completely monotone [13], i.e.,
@@
for all $t\in\mathbb{R}_{+}$ and $k\in\mathbb{Z}_{+}$. By Bernstein’s theorem, $g$ is completely monotone if and only if it is a conic combination of decaying exponentials, i.e., $g(t)=\sum_{i=1}^{n}G_{i}e^{-p_{i}t}$, where $G_{i}\geq 0$ and $p_{i}\geq 0$ for all $i\in\{1,\dots,n\}$. Relaxation systems are passive and reciprocal with only one type of energy storage. They admit a unique storage function that can be determined from the input-output behaviour of the system. Furthermore, every relaxation system has a state space realization that is internally of relaxation type [14], i.e., $B=C^{\ast}$ and $A\leq 0$.
@@
$(\mathcal{H}v)(t)=\int_{\mathbb{R}_{+}}g(t+\tau)v(\tau)\,\textrm{d}\tau$ (12)
The Hankel operator $\mathcal{H}$ maps past inputs to future outputs in the sense that the input $u(t)=v(-t)$ for all $t\leq 0$ and $u(t)=0$ for all $t>0$, results in the output $y(t)=(\mathcal{H}v)(t)$ for all $t\geq 0$. In [3], it is shown that the Hankel operator of a relaxation system is self-adjoint and nonnegative, hence cyclically monotone. This implies that $\mathcal{H}$ is the functional derivative of a quadratic (memory) functional, which serves as an intrinsic energy storage for the system.
Finally, we note that the results in this section are also valid for LTI systems with complex input, state and output spaces, which will be considered throughout this paper.
@@ IV LTSI systems and relaxation
In this section, we will define and characterize the notion of relaxation for a class of linear time-and-space-invariant (LTSI) systems. The definition of relaxation for LTSI systems will be a natural generalization of the definition for LTI systems, namely, it will be expressed as complete monotonicity of the (operator-valued) impulse response. The spatial invariance of the LTSI system will allow us to view it as a family of LTI systems via the Fourier transform. This will allow us to characterize relaxation of the LTSI system as relaxation of the corresponding family of LTI systems.
@@
$\mathcal{G}(t)=\mathcal{C}\exp\mathcal{A}(t)\mathcal{B}.$ (14)
@@
The fact that an LTSI system can be viewed as a family of LTI systems greatly simplifies its analysis. Indeed, many system theoretic properties of $\Sigma$ can be verified pointwise by verifying the analogous property for all $\hat{\Sigma}_{\omega}$, $\omega\in\mathbb{R}^{s}$, see, e.g., [16, 17]. This will be the overarching theme in this paper. With this in mind, as a natural generalization of the definition of relaxation for LTI systems, we define relaxation for LTSI systems as follows.
@@
for all $t>0$ and $k\in\mathbb{R}_{+}$. Since the Fourier transform is unitary, completely monotonicity of $\mathcal{G}$ is equivalent to complete monotonicity of $\hat{\mathcal{G}}$. The latter is a multiplication operator, hence its properties can be inferred from its symbol, i.e., the impulse responses $\hat{g}_{\omega}$, $\omega\in\mathbb{R}^{s}$. Consequently, relaxation can be verified pointwise, as shown below.
@@
As shown in [15, Theorem 2.1.13], the semigroup $\exp(\mathcal{A}t)$ is strongly differentiable and the $k$-th strong derivative is given by $\exp(\mathcal{A}t)\mathcal{A}^{k}$ for all $t>0$ on the dense domain $\mathscr{D}(\mathcal{A}^{k})\subset\mathcal{Z}$ Therefore, the impulse response $\mathcal{G}(t)$ is strongly differentiable with $k$-th strong derivative given by
@@
Since the Fourier transform is unitary, $\mathcal{A}\leq 0$ and $\mathcal{B}=\mathcal{C}^{\ast}$ if and only if $\hat{\mathcal{A}}\leq 0$ and ^$\hat{\mathcal{B}}=\hat{\mathcal{C}}^{\ast}$. Since the symbols $\hat{A}$, $\hat{B}$, and $\hat{C}$ of the multiplication operators $\hat{\mathcal{A}}$, $\hat{\mathcal{B}}$ and $\hat{\mathcal{C}}$ are continuous, $\hat{\mathcal{A}}\leq 0$ and ^$\hat{\mathcal{B}}=\hat{\mathcal{C}}^{\ast}$ if and only if $\hat{A}_{\omega}\leq 0$ and $\hat{B}_{\omega}=\hat{C}_{\omega}^{\ast}$ for all $\omega\in\mathbb{R}^{s}$. ∎
@@
Systems of the form (13) with $\mathcal{B}=\mathcal{C}^{\ast}$ are known as collocated systems [15, p.266]. The name reflects the practice of applying control and observation actions at the same point in distributed parameter systems. We refer to a collocated system of the form (13) with $\mathcal{A}=\mathcal{A}^{\ast}$ as internally symmetric. If, in addition, $\mathcal{A}\leq 0$, then $\mathcal{A}$ is the infinitesimal generator of a contraction semigroup, see [15, Section 2.3] for details. The latter is a mild form of stability which is implied by, e.g., exponential stability. Therefore, internal relaxation in LTSI systems is a combination of internal symmetry and stability, just like internal relaxation in LTI systems [6, 14].

### v2

@@ I Introduction
The remainder of this paper is organized as follows. In Section II, we introduce diffusion as a natural example of spatio-temporal relaxation. Section III contains notation and preliminaries. In Section IV, we define and characterize (internal) relaxation for LTSI systems. In Section V, we show that internal relaxation implies impedance passivity with storage determined from the external behaviour. In Section VI, we formalize the connection between internal and external by establishing properties of the Hankel operator of an LTSI relaxation system. Finally, Section VII summarizes our findings and discusses directions for further work.
@@ II Diffusion and relaxation
To motivate the content of this paper, we argue that diffusion is the physical phenomenon that should underline the system theoretic property of relaxation for LTSI systems. Historically, relaxation was first identified [8] as the system theoretic property of passive electrical circuits with only one type of energy storage element, i.e., RC or RL circuits. This characterization naturally extends to spatially distributed systems by considering an infinite transmission line composed of resistors and capacitors periodically arranged along a spatial axis. Figure 1 illustrates the schematic of such an infinite distributed RC transmission line. Taking the spatial discretization interval $\Delta x\rightarrow 0$, this discrete electrical network is precisely governed by the one-dimensional diffusion equation [9], i.e.
@@
where $V(t,x)$ and $I(t,x)$ are the voltage and current at time $t$ and position $x$. Denoting the variables $y(t,x)=V(t,x)$, $u(t,x)=I(t,x)/C$, and the coefficient $\alpha=1/RC$, the latter simplifies to the standard (heat) diffusion equation
@@
$y(t,x)=\int_{-\infty}^{t}\int_{-\infty}^{\infty}g(t-\tau,x-\xi)u(\tau,\xi)\,\textrm{d}\xi\textrm{d}\tau,$ (3)
@@
The heat kernel clearly exhibits the relaxation property when regarded as a spatially extended function of time. In particular, Figure 2(a) illustrates the spatial profile of (4) at various times $t$. It shows the relaxation property, as the spatial profile flattens monotonically towards equilibrium as $t$ increases. However, Figure 2(b) shows the kernel’s temporal behavior at a fixed location, highlighting that the kernel is not a completely monotone function in time, as one might expect from a classical LTI relaxation system.
@@
The diffusion process provides a clear physical connection from classical relaxation theory to its infinite-dimensional, spatially extended counterpart. This is further reinforced by the recent work [11], where it is explicitly demonstrated that the spatial discretization of the diffusion equation leads to a finite-dimensional LTI relaxation system. In this paper, we extend the theory of LTI relaxation systems to a class of LTSI system, with the diffusion process as a canonical example of an LTSI relaxation system.
xxt=1t=1t=2t=2t=5t=5t=20t=20
ttx=0x=0x=0.5x=0.5x=1x=1x=2x=2
@@ III Notation and preliminaries
We use mostly standard notation. We denote the sets of nonnegative real numbers and integers by $\mathbb{R}_{+}$ and $\mathbb{Z}_{+}$, respectively. We denote the inner-product on a (complex) Hilbert space $\mathcal{X}$ by $\langle\cdot,\cdot\rangle_{\mathcal{X}}$ and the induced norm by $\lVert\cdot\rVert_{\mathcal{X}}$. We omit the subscript if $\mathcal{X}=\mathbb{C}^{n}$ or there is no risk of ambiguity. We denote the Banach space of bounded linear operators from $\mathcal{X}$ to another Hilbert space $\mathcal{Y}$ by $\mathscr{B}(\mathcal{X},\mathcal{Y})$, and the operator norm by $\lVert\cdot\rVert$. We use the shorthand notation $\mathscr{B}(\mathcal{X},\mathcal{X})=\mathscr{B}(\mathcal{X})$. We also consider unbounded linear operators $\mathcal{A}:\mathscr{D}(\mathcal{A})\to\mathcal{Y}$, where $\mathscr{D}(\mathcal{A})\subset\mathcal{X}$ is the domain of $\mathcal{A}$. We say that $\mathcal{A}$ is densely defined if $\mathscr{D}(\mathcal{A})\subset\mathcal{X}$ is dense. If $\mathcal{A}$ is densely defined, then it has an adjoint $\mathcal{A}^{\ast}:\mathscr{D}(\mathcal{A}^{\ast})\to\mathcal{X}$, where $\mathcal{D}(\mathcal{A}^{\ast})\subset\mathcal{Y}$ is closed. We say that $\mathcal{A}$ is self-adjoint if $\mathscr{D}(\mathcal{A})=\mathscr{D}(\mathcal{A}^{\ast})$ and $\langle\mathcal{A}x,y\rangle=\langle x,\mathcal{A}y\rangle$ for all $x,y\in\mathscr{D}(\mathcal{A})$. We say that a self-adjoint $\mathcal{A}$ is nonnegative if $\langle\mathcal{A}x,x\rangle\geq 0$ for all $x\in\mathscr{D}(\mathcal{A})$, and we write $\mathcal{A}\geq 0$ to indicate that $\mathcal{A}$ is self-adjoint and nonnegative.
@@ III-A Square integrable functions and the Fourier transform
$\langle f,g\rangle_{L_{2}(X,\mathcal{X})}=\int_{X}\langle f(x),g(x)\rangle_{\mathcal{X}}\,\textrm{d}x.$ (7)
@@
$\hat{f}(\omega)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2\pi)^{-s/2}}\int_{\mathbb{R}^{s}}e^{-j\langle\omega,x\rangle}f(x)\,\textrm{d}x.$ (8)
@@
It is well-known [12] that the Fourier transform converges for absolutely integrable functions. Due to Plancherel’s theorem, the Fourier transform can be extended to a unitary operator $\mathcal{F}\in\mathscr{B}(L_{2}(\mathbb{R}^{s},\mathbb{C}^{n}))$, i.e., $\mathcal{F}^{-1}=\mathcal{F}^{\ast}$. Since $\mathcal{F}$ agrees with the Fourier transform on the dense subspace of absolutely integrable functions, we refer to $\mathcal{F}f$ as the Fourier transform of $f$, and we write $\hat{f}=\mathcal{F}f$.
@@
A key property of the Fourier transform that we exploit in this paper is that translation-invariant operators are transformed into multiplication operators. The Fourier transform of a densely defined operator $\mathcal{A}$ on $L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$ is the densely defined operator $\hat{\mathcal{A}}=\mathcal{F}\mathcal{A}\mathcal{F}^{\ast}$ on $L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$, such that $\hat{\mathcal{A}}\hat{f}\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$ is the Fourier transform of $\mathcal{A}f\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{n})$. For each $\bar{x}\in\mathbb{R}^{s}$, the translation operator $\mathcal{T}_{\bar{x}}\in\mathscr{B}(L_{2}(\mathbb{R}^{s},\mathbb{C}^{n}))$ is given by $(\mathcal{T}_{\bar{x}}f)(x)=f(x+\bar{x})$. A densely defined operator $\mathcal{A}$ is translation-invariant if $\mathcal{T}_{\bar{x}}:\mathscr{D}(\mathcal{A})\to\mathscr{D}(\mathcal{A})$ and $\mathcal{T}_{\bar{x}}\mathcal{A}=\mathcal{A}\mathcal{T}_{\bar{x}}$ for all $\bar{x}\in\mathbb{R}^{s}$. The Fourier transform of a translation-invariant operator is a multiplication operator [12, Section 3.6.7], i.e., there exists a function $\hat{A}:\mathbb{R}^{s}\to\mathbb{C}^{n\times n}$ such that $(\hat{\mathcal{A}}\hat{f})(\omega)=\hat{A}(\omega)\hat{f}(\omega)$. The function $\hat{A}$ is called the symbol of the operator $\hat{\mathcal{A}}$. It is well-known that the operators $\hat{\mathcal{A}}$ and, thus, $\mathcal{A}$ are bounded if and only if the symbol $\hat{A}$ is bounded.
@@
Throughout this paper, we consider functions that depend on a temporal variable $t$ and a spatial (frequency) variable $x$ ($\omega$). We denote the dependence on the spatial (frequency) variable with a subscript, e.g., we write $f_{x}$ instead of $f(x)$. In particular, we often consider functions of the form $f:\mathbb{T}\to L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$, where $\mathbb{T}$ is a temporal domain and $\mathbb{R}^{s}$ is a spatial (frequency) domain. We interpret such functions as mapping the temporal variable $t\in\mathbb{T}$ to a spatial (frequency) profile $f(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$. We denote the dependence on the spatial (frequency) variable with a subscript, i.e., we write $f_{x}(t)=f(t)(x)$, so that the function $f_{x}:\mathbb{T}\to\mathbb{C}^{n}$ is the temporal trajectory at a given point in space $x\in\mathbb{R}^{s}$. Finally, we write $\hat{f}$ to denote the spatial Fourier transform of $f$, i.e., the function $\hat{f}:\mathbb{T}\to L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ such that $\hat{f}(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ is the Fourier transform of $f(t)\in L_{2}(\mathbb{R}^{s},\mathbb{C}^{m})$ for all $t\in\mathbb{T}$.
@@ III-B LTI relaxation systems
$\Sigma:\left\{\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}z(t)&=Az(t)+Bu(t),\\ y(t)&=Cz(t),\end{aligned}\right$. (9)
@@
$y(t)=\int_{-\infty}^{t}g(t-\tau)u(\tau)\,\textrm{d}\tau.$ (10)
The system $\Sigma$ is of relaxation type [6] if $m=p$ and its impulse response $g$ is devoid of any oscillatory behaviour, expressed as $g$ being completely monotone [13, Chapter IV], i.e.,
@@
for all $t\in\mathbb{R}_{+}$ and $k\in\mathbb{Z}_{+}$. By Bernstein’s theorem [14], $g$ is completely monotone if and only if it is a conic combination of decaying exponentials, i.e., $g(t)=\sum_{i=1}^{n}G_{i}e^{-p_{i}t}$, where $G_{i}\geq 0$ and $p_{i}\geq 0$ for all $i\in\{1,\dots,n\}$. Relaxation systems are passive and reciprocal with only one type of energy storage. They admit a unique storage function that can be determined from the input-output behaviour of the system. Furthermore, every relaxation system has a state space realization that is internally of relaxation type [15], i.e., $B=C^{\ast}$ and $A\leq 0$.
@@
$(\mathcal{H}v)(t)=\int_{\mathbb{R}_{+}}g(t+\tau)v(\tau)\,\textrm{d}\tau.$ (12)
The Hankel operator $\mathcal{H}$ maps past inputs to future outputs in the sense that the input u(t)=v(t)u(t)=v(-t) for all $t\leq 0$ and $u(t)=0$ for all $t>0$, results in the output $y(t)=(\mathcal{H}v)(t)$ for all $t\geq 0$. In [3], it is shown that the Hankel operator of a relaxation system is self-adjoint and nonnegative, hence cyclically monotone. This implies that $\mathcal{H}$ is the functional derivative of a quadratic (memory) functional, which serves as an intrinsic energy storage for the system.
Finally, we note that the results in this section are also valid for LTI systems with complex input, state and output spaces, which are considered throughout this paper.
@@ IV LTSI systems and relaxation
In this section, we define and characterize the notion of relaxation for a class of linear time-and-space-invariant (LTSI) systems. The definition of relaxation for LTSI systems is a natural generalization of the definition for LTI systems, namely, it is expressed as complete monotonicity of the (operator-valued) impulse response. The spatial invariance of the LTSI system allows us to view it as a family of LTI systems via the Fourier transform. This allows us to characterize relaxation of the LTSI system as relaxation of the corresponding family of LTI systems.
@@
$\mathcal{G}(t)=\mathcal{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp(\mathcal{A}t)}\mathcal{B}.$ (14)
@@
The fact that an LTSI system can be viewed as a family of LTI systems greatly simplifies its analysis. Indeed, many system theoretic properties of $\Sigma$ can be verified pointwise by verifying the analogous property for all $\hat{\Sigma}_{\omega}$, $\omega\in\mathbb{R}^{s}$, see, e.g., [17, 18]. This is the overarching theme in this paper. With this in mind, as a natural generalization of the definition of relaxation for LTI systems, we define relaxation for LTSI systems as follows.
@@
for all $t>0$ and $k\in\mathbb{R}_{+}$. Since the Fourier transform is unitary, completely monotonicity of $\mathcal{G}$ is equivalent to complete monotonicity of $\hat{\mathcal{G}}$. The latter is a multiplication operator, hence its properties can be inferred from its symbol, i.e., the impulse responses $\hat{g}_{\omega}$, $\omega\in\mathbb{R}^{s}$. Consequently, relaxation can be verified pointwise, as shown in the following theorem.
@@
As shown in [16, Theorem 2.1.13], the semigroup $\exp(\mathcal{A}t)$ is strongly differentiable and the $k$-th strong derivative is given by $\exp(\mathcal{A}t)\mathcal{A}^{k}$ for all $t>0$ on the dense domain $\mathscr{D}(\mathcal{A}^{k})\subset\mathcal{Z}$ . Therefore, the impulse response $\mathcal{G}(t)$ is strongly differentiable with $k$-th strong derivative given by
@@
Since the Fourier transform is unitary, $\mathcal{A}\leq 0$ and $\mathcal{B}=\mathcal{C}^{\ast}$ if and only if $\hat{\mathcal{A}}\leq 0$ and $\hat{\mathcal{B}}=\hat{\mathcal{C}}^{\ast}$. Since the symbols $\hat{A}$, $\hat{B}$, and $\hat{C}$ of the multiplication operators $\hat{\mathcal{A}}$, $\hat{\mathcal{B}}$ and $\hat{\mathcal{C}}$ are continuous, $\hat{\mathcal{A}}\leq 0$ and $\hat{\mathcal{B}}=\hat{\mathcal{C}}^{\ast}$ if and only if $\hat{A}_{\omega}\leq 0$ and $\hat{B}_{\omega}=\hat{C}_{\omega}^{\ast}$ for all $\omega\in\mathbb{R}^{s}$. ∎
@@
Systems of the form (13) with $\mathcal{B}=\mathcal{C}^{\ast}$ are known as collocated systems [16, p.266]. The name reflects the practice of applying control and observation actions at the same point in distributed parameter systems. We refer to a collocated system of the form (13) with $\mathcal{A}=\mathcal{A}^{\ast}$ as internally symmetric. If, in addition, $\mathcal{A}\leq 0$, then $\mathcal{A}$ is the infinitesimal generator of a contraction semigroup, see [16, Section 2.3] for details. The latter is a mild form of stability which is implied by, e.g., exponential stability. Therefore, internal relaxation in LTSI systems is a combination of internal symmetry and stability, just like internal relaxation in LTI systems [6, 15].